If α is in first quadrant= Π/19.Find value of sin 23α-sin 3α/sin 16α+sin 4α

Grade:11

1 Answers

Akshay
8 years ago
Hint : use [ sin x – sin y = 2* sin{(x-y)/2} * cos{(x+y)/2} ] and [ sin x + sin y = 2* sin{(x+y)/2} * cos{(x-y)/2} ], 
and then use cos x = – cos(pi-x)